# Kuznecov formulae for fractal measures

### v1

@@ Abstract.
Assume that $\mu$ is $s$-Ahlfors regular for some $s\in(0,n)$ and admits an averaged $s$-density constant $A_{\mu}$. We prove that
@@
The hypotheses of $s$-Ahlfors regularity and the averaged $s$-density condition are essentially optimal for such a one-term asymptotic, and in general the remainder $o(\lambda^{n-s})$ cannot be improved uniformly to a power-saving error term. This extends the classical Kuznecov formulae of Zelditch [Zel92] for smooth submanifold measures to a broad class of singular and fractal measures.
@@ 1. Introduction
where $B^{n-k}\subset\mathbb{R}^{n-k}$ is the Euclidean unit ball and $\operatorname{vol}(H)$ denotes the kk–dimensional Riemannian volume of $H$. Thus the leading growth of the Kuznecov sum is governed by the ambient dimension $n$, the submanifold dimension $k$, and an explicit geometric factor.
@@
Throughout the paper we use the following convention for Euclidean ball volumes. For $m>0$ (not necessarily an integer), we write $\operatorname{vol}(B^{m})$ for the “mm–dimensional volume” of the Euclidean unit ball in formal dimension $m$, defined by
$\operatorname{vol}(B^{m}):=\frac{\pi^{m/2}}{\Gamma\bigl(\frac{m}{2}+1\bigr)}.$
When $m\in\mathbb{N}$, this coincides with the usual Lebesgue mm–volume of the unit ball in $\mathbb{R}^{m}$. For an embedded kk–dimensional submanifold $H\subset M$ we write $\operatorname{vol}(H)$ for its kk–dimensional Riemannian volume.
We now describe the small-scale geometric hypotheses on $\mu$ that will play the role of “dimension” and “density” in our setting. Throughout, $B(x,r)$ denotes the geodesic ball of radius $r$ centered at $x\in M$ with respect to the Riemannian distance $d_{g}$.
Definition 1.1 ($s$-Ahlfors regular measure).
Let $s\in(0,n)$. A finite Borel measure $\mu$ on $M$ is called $s$-Ahlfors regular if there exist constants $c,C>0$ and $r_{0}>0$ such that
This condition is standard in geometric measure theory and implies that the Hausdorff dimension of $\mu$ is $s$. See, for example, [Mat15]. It is satisfied by Riemannian volume measures on $s$-dimensional embedded submanifolds, as well as by many self-similar and self-conformal measures on fractals.
To obtain a one-term asymptotic for (1.1), one needs a stabilization of the averaged small-scale behavior of $\mu$. For $r\geq 0$ define the distance distribution
@@
Let $s\in(0,n)$. We say that $\mu$ admits an averaged $s$-density if there exists a constant $A_{\mu}\in(0,\infty)$ (called the averaged $s$-density constant of $\mu$) such that
@@
We can now state our main result. It shows that under the Ahlfors regularity assumption and the averaged $s$-density condition (1.5), the Kuznecov sum (1.1) has a one-term asymptotic with exponent nsn-s and an explicit leading constant depending only on $n$ and $s$.
@@
Let $(M,g)$ be a compact, connected Riemannian manifold of dimension $n\geq 2$. Fix $s\in(0,n)$ and let $\mu$ be a finite Borel measure on $M$ which is $s$-Ahlfors regular in the sense of (1.3) and admits an averaged $s$-density constant $A_{\mu}$ in the sense of (1.5). Then the Kuznecov sum (1.1) satisfies
@@
The proof combines a heat-kernel regularization of the Kuznecov sum with a Tauberian argument, with the leading term determined by the small-scale distance distribution of $\mu$.
@@
The averaged $s$-density condition (1.5) is automatically satisfied under the stronger pointwise density hypothesis. Namely, suppose that $\mu$ is $s$-Ahlfors regular and admits a pointwise $s$-density $\theta$ in the sense that
@@
exists for $\mu$-a.e. $x$. Then by the Ahlfors upper bound $\mu(B(x,r))/(\operatorname{vol}(B^{s})r^{s})\lesssim 1$ and dominated convergence,
$\frac{F(r)}{\operatorname{vol}(B^{s})\,r^{s}}=\int_{M}\frac{\mu(B(x,r))}{\operatorname{vol}(B^{s})\,r^{s}}\,d\mu(x)\longrightarrow\int_{M}\theta(x)\,d\mu(x),\qquad(r\to 0^{+}).$
Hence (1.5) holds with
$A_{\mu}=\int_{M}\theta\,d\mu.$
In particular, (1.5) is strictly weaker than the existence of a pointwise $s$-density.
The leading exponent in (1.7) is nsn-s, determined by the ambient dimension $n$ and the Hausdorff dimension $s$ of $\mu$, and the leading coefficient factors as a universal constant times the averaged $s$-density constant $A_{\mu}$. In the special case where $\mu=dV_{H}$ is the Riemannian volume measure on an embedded submanifold $H$ of dimension $s$, $\mu$ admits a pointwise $s$-density $\theta$ and hence satisfies the averaged $s$-density condition (1.5) by Remark 1.4. In fact $\theta\equiv 1$ and $A_{\mu}=\operatorname{vol}(H)$, so (1.7) reduces to
@@
whose leading term agrees with Zelditch’s Kuznecov formulae (1.2) with $k=s$. Thus Theorem 1.3 extends the classical Kuznecov formulae to a large class of singular and fractal measures. To the best of our knowledge, Theorem 1.3 is the first Kuznecov-type asymptotic for fractal measures.
Moreover, as discussed in [WX23a], the left-hand side of Zelditch’s Kuznecov formula (1.2) can be viewed as “audible” spectral information associated with the submanifold $H$, while the right-hand side encodes geometric information about $H$, namely its dimension $k$ and its $k$-dimensional Riemannian volume $\operatorname{vol}(H)$. Theorem 1.3 extends this audible–geometric correspondence from smooth submanifold measures to $s$-Ahlfors regular measures admitting an averaged $s$-density constant, so that the Hausdorff dimension $s$ and the effective “volume” $A_{\mu}$ of the measure are also audible through the leading asymptotics of $N_{\mu}(\lambda)$.
@@
denote the Riesz energy of $\mu$ with respect to the Riemannian distance $d_{g}(x,y)$ on $M$. There is an earlier work of Hare and Roginskaya [HR03] which is closely related in spirit to our results. They established energy formulae on compact Riemannian manifolds which relate $I_{u}(\mu)$ to weighted sums of the projections of $\mu$ onto Laplace eigenspaces. In Section 4 we show that their identities imply the following general upper bound without assuming the existence of an averaged $s$-density constant.
@@
In particular, if $\mu$ is $s$-Ahlfors regular for some $s\in(0,n)$, then $I_{u}(\mu)<\infty$ for every $u<s$, and hence for each $\varepsilon>0$ there exists a constant $C_{2}=C_{2}(M,g,s,\varepsilon,\mu)$ such that
@@
We point out that the assumptions in Theorem 1.3 and Theorem 1.5 are essentially optimal. In Section 5 we give examples showing, on the one hand, that one cannot dispense with the $s$-Ahlfors regularity assumption if one wishes to control the exponent purely in terms of the Hausdorff dimension of $\mu$, and on the other hand, that some stabilization of the small-scale distance distribution—such as the averaged $s$-density condition (1.5)—is needed to obtain a genuine one-term asymptotic. Under $s$-Ahlfors regularity alone one can at best expect upper and lower bounds of order $\lambda^{n-s}$, but not the existence of a limit for $N_{\mu}(\lambda)/\lambda^{n-s}$.
In addition to these structural restrictions, we show in Section 5 that the remainder term in Theorem 1.3 cannot, in general, be improved to a uniform power saving over the class of measures under consideration. More precisely, we construct explicit examples showing that there is no exponent $\delta>0$ such that for every $s$-Ahlfors regular measure with an averaged $s$-density constant $A_{\mu}$ one has
$N_{\mu}(\lambda)=C_{n,s}A_{\mu}\,\lambda^{n-s}+O\bigl(\lambda^{n-s-\delta}\bigr),$
with an implicit constant depending only on $(M,g,n,s,\mu)$. The $o(\lambda^{n-s})$ remainder in (1.7) is therefore essentially sharp at this level of generality.
@@
In Section 2 we introduce a heat-kernel regularization of the Kuznecov sum, and establish its small-time asymptotics in terms of the averaged $s$-density constant $A_{\mu}$. Section 3 interprets this smoothed Kuznecov sum as a Laplace–Stieltjes transform and applies a Tauberian theorem to prove Theorem 1.3 and identify the constant $C_{n,s}$. In Section 4 we revisit the energy identities of Hare and Roginskaya to derive general polynomial upper bounds for $N_{\mu}(\lambda)$, proving Theorem 1.5. Finally, Section 5 is devoted to sharpness: we give examples showing that the Ahlfors regularity and averaged $s$-density hypotheses in Theorem 1.3 are essentially optimal, and that the $o(\lambda^{n-s})$ remainder cannot, in general, be improved to a uniform power-saving error term.
Acknowledgement
This project is supported by the National Key R&D Program of China under Grant No. 2022YFA1007200, the Natural Science Foundation of China under Grant No. 12571107, and the Zhejiang Provincial Natural Science Foundation of China under Grant No. LR25A010001. The author thanks Xucheng Hu and Diankun Liu for carefully reading an earlier draft of this paper.
@@ 2.1. Heat kernel asymptotics near the diagonal
There exist $r_{0}>0$, $t_{0}>0$, and smooth functions $u_{k}(x,y)$ on $\{(x,y)\in M\times M:d_{g}(x,y)<2r_{0}\}$, $k=0,1,2,\dots$, with $u_{0}(x,x)\equiv 1$, such that for every $N\in\mathbb{N}$ we have
@@ 2.3. A Gaussian average
Assume that $\mu$ satisfies (1.3) and (1.5). Then, as $t\to 0^{+}$,
@@
with $\mathcal{E}(r)=o(r^{s})$ as $r\to 0^{+}$. Fix $r_{0}>0$ as in (1.3). Split
@@
On $\delta,r_{0}]$ we use the upper Ahlfors bound $F(r)\leq Cr^{s}$ to deduce that $|\mathcal{E}(r)|\leq C^{\prime}r^{s}$ on this interval. Thus
@@ 2.4. Small-time asymptotics of the smoothed Kuznecov sum
Assume that $\mu$ satisfies (1.3) and (1.5). Then
@@
For the second integral we use (2.3) and the Ahlfors bound via $F(r)$:
@@ 3.2. A Tauberian theorem for Laplace–Stieltjes transforms
$(2\pi)^{-(n-s)}\,\operatorname{vol}(B^{n-s})=2^{-(n-s)}\pi^{-(n-s)}\cdot\frac{\pi^{(n-s)/2}}{\Gamma\bigl(\frac{n-s}{2}+1\bigr)}=2^{-(n-s)}\pi^{-(n-s)/2}\frac{1}{\Gamma\bigl(\frac{n-s}{2}+1\bigr)}=C_{n,s}.$
@@
Remark 4.2.
Estimates of the form (4.1) are proved in [HR03] by expressing suitable powers of the Laplace–Beltrami operator as pseudodifferential operators and identifying their kernels with Riesz-type potentials. In particular, $I_{u}(\mu)$ is comparable, up to multiplicative constants depending only on $(M,g,n,u)$, to the quadratic form associated with the operator $(I-\Delta_{g})^{-(n-u)/2}$ evaluated on $\mu$. See [HR03] for details.
@@
To obtain the “in particular” statement, it remains to note that if $\mu$ is $s$-Ahlfors regular for some $s\in(0,n)$, then $I_{u}(\mu)<\infty$ for every $u<s$. This is a standard estimate, but we include a brief argument in our setting for completeness.
@@
Let $\mu$ be an $s$-Ahlfors regular measure on $M$ with $s\in(0,n)$, and let $u\in(0,s)$. Then $I_{u}(\mu)<\infty$.
@@
$I_{u}(\mu)=\int_{0}^{\operatorname{diam}(M)}ur^{-u-1}F(r)\,dr+\mu(M)^{2}\,\operatorname{diam}(M)^{-u},$
@@
$u\int_{0}^{r_{0}}r^{-u-1}F(r)\,dr\leq uC\int_{0}^{r_{0}}r^{s-u-1}\,dr=\frac{uC}{s-u}\,r_{0}^{\,s-u}<\infty,$
since su>0s-u>0. On $r_{0},\operatorname{diam}(M)]$,
@@
Theorem 1.5 shows that for each fixed $u<s$ one can control the growth of the Kuznecov sum by $\lambda^{n-u}$ under the mild hypothesis $I_{u}(\mu)<\infty$, and in particular for $s$-Ahlfors regular measures. Our main theorem, Theorem 1.3, goes further by identifying the sharp exponent nsn-s and the exact leading constant in the asymptotic (1.7), under the additional assumption of the averaged $s$-density condition. As discussed in the introduction, the methods of [HR03] treat $u$ as a fixed parameter and do not provide the uniform control in $u$ near $s$ that is needed to pass to the endpoint $u=s$. This is why we turn to a more refined analysis based on the heat kernel and Tauberian theory.
5. Sharpness of Theorem 1.3
In this section, we discuss the sharpness of Theorem 1.3. We explain the necessity of the two structural hypotheses, namely $s$-Ahlfors regularity and the averaged $s$-density condition (1.5), for obtaining the exponent nsn-s and a genuine one-term asymptotic. We also show that the remainder $o(\lambda^{n-s})$ in (1.7) is essentially optimal at this level of generality: in general it cannot be improved uniformly to a power-saving error term over the class of measures under consideration.
5.1. Necessity of the Ahlfors regularity assumption
It is natural to ask whether the exponent nsn-s in Theorem 1.3 (or in the upper bound of Theorem 1.5) could be expressed solely in terms of the Hausdorff dimension of $\mu$, without any Ahlfors regularity hypothesis. This is not possible in general.
Let $H_{1},H_{2}\subset M$ be disjoint embedded submanifolds with dimensions $s_{1}<s_{2}$, and let $\mu_{i}=dV_{H_{i}}$ and $\mu=\mu_{1}+\mu_{2}$. Then
$\dim_{H}(\mu)=\max(s_{1},s_{2})=s_{2},$
but the growth of $N_{\mu}(\lambda)$ is governed by the smaller dimension $s_{1}$.
To see this, write
$a_{j}:=\int_{H_{1}}e_{j}\,dV_{H_{1}},\qquad b_{j}:=\int_{H_{2}}e_{j}\,dV_{H_{2}}.$
Then
$N_{\mu}(\lambda)=\sum_{\lambda_{j}\leq\lambda}|a_{j}+b_{j}|^{2}=\sum_{\lambda_{j}\leq\lambda}|a_{j}|^{2}+\sum_{\lambda_{j}\leq\lambda}|b_{j}|^{2}+2\Re\sum_{\lambda_{j}\leq\lambda}a_{j}\overline{b_{j}}.$
By Cauchy–Schwarz,
$\Bigl|\sum_{\lambda_{j}\leq\lambda}a_{j}\overline{b_{j}}\Bigr|\leq\Bigl(\sum_{\lambda_{j}\leq\lambda}|a_{j}|^{2}\Bigr)^{1/2}\Bigl(\sum_{\lambda_{j}\leq\lambda}|b_{j}|^{2}\Bigr)^{1/2}.$
Hence we have the robust comparison
(5.1) $\bigl|N_{\mu}(\lambda)-N_{\mu_{1}}(\lambda)-N_{\mu_{2}}(\lambda)\bigr|\leq 2\,\bigl(N_{\mu_{1}}(\lambda)N_{\mu_{2}}(\lambda)\bigr)^{1/2}.$
By the smooth submanifold Kuznecov formulae,
$N_{\mu_{i}}(\lambda)\sim C_{n,s_{i}}\,\operatorname{vol}(H_{i})\,\lambda^{n-s_{i}}\qquad(i=1,2),$
so $N_{\mu_{2}}(\lambda)=o(N_{\mu_{1}}(\lambda))$ and the right-hand side of (5.1) is
$O\Bigl(\lambda^{\frac{n-s_{1}}{2}}\,\lambda^{\frac{n-s_{2}}{2}}\Bigr)=O\bigl(\lambda^{\,n-\frac{s_{1}+s_{2}}{2}}\bigr)=o\bigl(\lambda^{n-s_{1}}\bigr).$
Therefore $N_{\mu}(\lambda)=N_{\mu_{1}}(\lambda)+o(\lambda^{n-s_{1}})$, and in particular
$N_{\mu}(\lambda)\sim C_{n,s_{1}}\,\operatorname{vol}(H_{1})\,\lambda^{n-s_{1}}\qquad(\lambda\to\infty),$
even though $\dim_{H}(\mu)=s_{2}$. Thus, no exponent depending only on $\dim_{H}(\mu)$ can govern $N_{\mu}(\lambda)$ for arbitrary finite Borel measures.

### v2

@@ Abstract.
Assume that $\mu$ admits an averaged $s$-density constant $A_{\mu}$ with correlation dimension $s\in(0,n)$. We prove that
@@
The averaged $s$-density condition is necessary for such a one-term asymptotic, and in general, the remainder $o(\lambda^{n-s})$ is sharp in the sense that it cannot be improved uniformly to a power-saving error term. This extends the classical Kuznecov formula of Zelditch [Zel92] for smooth submanifold measures to a broad class of singular and fractal measures.
@@ 1. Introduction
where $B^{n-k}\subset\mathbb{R}^{n-k}$ is the Euclidean unit ball and $\operatorname{vol}(H)$ denotes the $k$-dimensional Riemannian volume of $H$. Thus the leading growth of the Kuznecov sum is governed by the ambient dimension $n$, the submanifold dimension $k$, and an explicit geometric factor.
@@
Throughout the paper we use the following convention for Euclidean ball volumes. For $s\in\mathbb{R}^{+}$, we write $\operatorname{vol}(B^{s})$ for the “$s$-dimensional volume” of the Euclidean unit ball in formal dimension $s$, defined by
$\operatorname{vol}(B^{s}):=\frac{\pi^{s/2}}{\Gamma\bigl(\frac{s}{2}+1\bigr)}.$
When $s\in\mathbb{Z}^{+}$, this coincides with the usual Lebesgue $s$-volume of the unit ball in $\mathbb{R}^{s}$. For an embedded $k$-dimensional submanifold $H\subset M$ we write $\operatorname{vol}(H)$ for its $k$-dimensional Riemannian volume.
We now describe the small-scale hypotheses on $\mu$ that will play the role of “dimension” and “volume” in our setting. Throughout, $B(x,r)$ denotes the geodesic ball of radius $r$ centered at $x\in M$ with respect to the Riemannian distance $d_{g}$.
For $r\geq 0$ define the distance distribution
@@
The small-scale growth rate of $F(r)$ encodes the pair statistics of $\mu$ and leads naturally to the correlation dimension of $\mu$, see for instance [MMR00].
Definition 1.1 (Correlation dimension).
Let $\mu$ be a finite Borel measure on $M$. If the following limit exists, we define the correlation dimension of $\mu$ to be
(1.4) $\dim_{C}(\mu):=\lim_{r\to 0^{+}}\frac{\log F(r)}{\log r}.$
To obtain a one-term asymptotic for (1.1), it is necessary to assume that the small-scale behavior of $F(r)$ stabilizes in the following sense.
@@
Let $s\in(0,n)$. We say that $\mu$ admits an averaged $s$-density constant $A_{\mu}\in(0,\infty)$ if
@@
When (1.5) holds, it forces $\dim_{C}(\mu)=s$, and $A_{\mu}$ is the corresponding averaged $s$-dimensional density constant arising from the distance distribution.
We can now state our main result. It shows that once the distance distribution $F(r)$ admits a stabilized small-scale asymptotic in the sense of (1.5), the Kuznecov sum (1.1) has a one-term asymptotic with exponent nsn-s and an explicit leading constant.
@@
Let $(M,g)$ be a compact, connected Riemannian manifold of dimension $n\geq 2$. Fix $s\in(0,n)$ and let $\mu$ be a finite Borel measure on $M$ admitting an averaged $s$-density constant $A_{\mu}$ in the sense of (1.5). Then
@@
The proof combines a heat-kernel regularization of the Kuznecov sum with a Tauberian argument, with the leading term determined by the small-scale distance distribution of $\mu$. Combining (1.7) with Weyl’s law for the eigenvalue counting function shows that the Fourier coefficients $\int_{M}e_{j}\,d\mu$ have a quantitative mean-square decay on average. The precise statement is recorded in the following corollary.
Corollary 1.4 (Average Fourier decay).
Under the hypotheses of Theorem 1.3, we have
$\frac{1}{\#\{j:\lambda_{j}\leq\lambda\}}\sum_{\lambda_{j}\leq\lambda}\Bigl|\int_{M}e_{j}\,d\mu\Bigr|^{2}\approx\lambda^{-s},\qquad(\lambda\to\infty).$
If one wishes to interpret the exponent $s$ more geometrically, a natural setting is the class of Ahlfors–David regular measures. Recall that a finite Borel measure $\mu$ on $M$ is $s$-Ahlfors–David regular if there exist constants $c,C>0$ and $r_{0}>0$ such that
For such measures it is standard that the Hausdorff dimension of $\mu$ equals $s$. Thus, within the Ahlfors–David regular class, the exponent in (1.7) may be written as $n-\dim_{H}(\mu)$.
@@
The averaged $s$-density condition (1.5) is automatically satisfied under a stronger pointwise density hypothesis, together with a uniform upper $s$-ball condition. Suppose that there exist $C>0$ and $r_{0}>0$ such that $\mu(B(x,r))\leq Cr^{s}$ for all $x\in\operatorname{supp}\mu$ and $0<r\leq r_{0}$, and that the pointwise $s$-density
@@
exists for $\mu$-a.e. $x$. Then dominated convergence implies
$\frac{F(r)}{\operatorname{vol}(B^{s})\,r^{s}}=\int_{M}\frac{\mu(B(x,r))}{\operatorname{vol}(B^{s})\,r^{s}}\,d\mu(x)\longrightarrow\int_{M}\theta(x)\,d\mu(x),\qquad(r\to 0^{+}),$
so (1.5) holds with $A_{\mu}=\int_{M}\theta\,d\mu$. In particular, (1.5) is strictly weaker than the existence of a pointwise $s$-density. Indeed, by a theorem of Preiss [Pre87], if a (Radon) measure has an $s$-density $\theta$ with $0<\theta(x)<\infty$ for $\mu$-a.e. $x$, then necessarily $s\in\mathbb{N}$ and $\mu$ is $s$-rectifiable. This integer-dimensional rigidity does not apply to the averaged density condition (1.5), which allows genuinely non-integer $s$. See Proposition 5.3.
In the special case where $\mu=dV_{H}$ is the Riemannian volume measure on an embedded submanifold $H\subset M$ of dimension $s$, $\mu$ admits a pointwise $s$-density $\theta$ and hence satisfies the averaged $s$-density condition (1.5) by Remark 1.5. In fact $\theta\equiv 1$ and $A_{\mu}=\operatorname{vol}(H)$, so (1.7) reduces to
@@
whose leading term agrees with Zelditch’s Kuznecov formula (1.2) with $k=s$.
More generally, Theorem 1.3 implies that the same conclusion holds when $H$ has only weak regularity. For instance, suppose that $H\subset M$ is a compact embedded Lipschitz $s$-dimensional submanifold with Lipschitz boundary. Let $\mu$ be the $s$-dimensional Hausdorff measure induced by $d_{g}$ on $H$. Then $\mu$ admits a pointwise $s$-density $\theta\equiv 1$ for $\mu$-a.e. $x\in H$. Hence $\mu$ satisfies (1.5) and Theorem 1.3 yields the same leading asymptotic with $A_{\mu}=\operatorname{vol}(H)$. This provides a Kuznecov asymptotic at Lipschitz regularity, whereas the classical results for submanifold measures assume that $H$ is smooth and boundaryless.
In summary, Theorem 1.3 extends the classical Kuznecov formulae to a broad class of singular and fractal measures. To the best of our knowledge, Theorem 1.3 provides the first Kuznecov-type asymptotic for fractal measures.
Moreover, as discussed in [WX23a], the left-hand side of Zelditch’s Kuznecov formula (1.2) can be viewed as “audible” spectral information associated with the submanifold $H$, while the right-hand side encodes geometric information about $H$, namely its dimension $k$ and its $k$-dimensional Riemannian volume $\operatorname{vol}(H)$. Theorem 1.3 extends this audible–geometric correspondence to measures $\mu$ whose distance distribution admits an averaged $s$-density constant $A_{\mu}$ in the sense of (1.5). In this case the correlation dimension $s$ and the effective “volume” $A_{\mu}$ are audible through the leading asymptotics of $N_{\mu}(\lambda)$. If, in addition, $\mu$ is $s$-Ahlfors–David regular, then its Hausdorff dimension is also audible.
@@
denote the Riesz energy of $\mu$ with respect to the Riemannian distance $d_{g}(x,y)$ on $M$. There is an earlier work of Hare and Roginskaya [HR03] which is closely related in spirit to our results. They established energy formulae on compact Riemannian manifolds which relate $I_{u}(\mu)$ to weighted sums of the projections of $\mu$ onto Laplace eigenspaces. In Section 4 we show that their identities imply the following general upper bound, without assuming the existence of an averaged $s$-density constant.
@@
In particular, if $\mu$ has correlation dimension $\dim_{C}(\mu)=s$ for some $s\in(0,n)$, then $I_{u}(\mu)<\infty$ for every $u<s$, and hence for each $\varepsilon>0$ there exists a constant $C_{2}=C_{2}(M,g,s,\varepsilon,\mu)$ such that
@@
We record two statements, proved in Section 5, which together explain why the averaged $s$-density hypothesis (1.5) in Theorem 1.3 is essentially optimal for a one-term asymptotic, and why the remainder $o(\lambda^{n-s})$ in (1.7) cannot be improved uniformly.
Even for Ahlfors–David regular measures, a genuine one-term asymptotic requires stabilization of the small-scale geometry of $\mu$. The next proposition shows that without (1.5), a one-term asymptotic need not hold.
Proposition 1.7 (Necessity of averaged $s$-density).
There exists an $s$-Ahlfors–David regular measure $\mu$ for which the quantity $N_{\mu}(\lambda)/\lambda^{n-s}$ fails to converge as $\lambda\to\infty$.
Beyond these structural issues, we also show that the remainder $o(\lambda^{n-s})$ in (1.7) is essentially optimal at this level of generality: there is no uniform power-saving improvement valid even over the Ahlfors–David regular subclass.
Proposition 1.8 (Sharpness of the remainder).
Fix $\delta\in(0,1)$. There exists an $s$-Ahlfors–David regular measure $\mu$ admitting an averaged $s$-density constant $A_{\mu}$ such that
$N_{\mu}(\lambda)=C_{n,s}A_{\mu}\,\lambda^{n-s}+O(\lambda^{n-s-\delta})$
fails to hold.
@@
In Section 2 we introduce a heat-kernel regularization of the Kuznecov sum and establish its small-time asymptotics in terms of the averaged $s$-density constant $A_{\mu}$. Section 3 interprets this smoothed Kuznecov sum as a Laplace–Stieltjes transform and applies a Tauberian theorem to prove Theorem 1.3 and identify the constant $C_{n,s}$. In Section 4 we revisit the energy identities of Hare and Roginskaya to derive general polynomial upper bounds for $N_{\mu}(\lambda)$, proving Theorem 1.6. Finally, Section 5 is devoted to sharpness and examples: we give examples showing that the averaged $s$-density hypothesis in Theorem 1.3 is essentially optimal for a genuine one-term asymptotic, and that the $o(\lambda^{n-s})$ remainder cannot, in general, be improved uniformly to a power-saving error term.
Acknowledgements
This project is supported by the National Key R&D Program of China under Grant No. 2022YFA1007200, the Natural Science Foundation of China under Grant No. 12571107, and the Zhejiang Provincial Natural Science Foundation of China under Grant No. LR25A010001. The author thanks Shi-Lei Kong for some helpful discussions. The author thanks Xucheng Hu and Diankun Liu for carefully reading an earlier draft of this paper.
@@ 2.1. Heat kernel asymptotics near the diagonal
There exist $r_{0}>0$, $t_{0}>0$, and smooth functions $u_{k}(x,y)$ on $\{(x,y)\in M\times M:d_{g}(x,y)<2r_{0}\}$, $k=0,1,2,\dots$, with $u_{0}(x,x)\equiv 1$, such that for every $N\in\mathbb{Z}^{+}$ we have
@@ 2.3. A Gaussian average
Assume that $\mu$ satisfies (1.5). Then, as $t\to 0^{+}$,
@@
with $\mathcal{E}(r)=o(r^{s})$ as $r\to 0^{+}$. Fix $r_{0}>0$ to be a small constant. Split
@@
On $\delta,r_{0}]$ we use the upper bound $F(r)\leq Cr^{s}$ to deduce that $|\mathcal{E}(r)|\leq C^{\prime}r^{s}$ on this interval. Thus
@@ 2.4. Small-time asymptotics of the smoothed Kuznecov sum
Assume that $\mu$ satisfies (1.5). Then
@@
For the second integral we use (2.3):
@@ 3.2. A Tauberian theorem for Laplace–Stieltjes transforms
$(2\pi)^{-(n-s)}\,\operatorname{vol}(B^{n-s})=2^{-(n-s)}\pi^{-(n-s)}\cdot\frac{\pi^{(n-s)/2}}{\Gamma\bigl(\frac{n-s}{2}+1\bigr)}=C_{n,s}.$
@@
To obtain the second statement of Theorem 1.6, it remains to note that if $\mu$ has correlation dimension $\dim_{C}(\mu)=s$ for some $s\in(0,n)$, then $I_{u}(\mu)<\infty$ for every $u<s$. This is a standard estimate, and is the reason why the correlation dimension is sometimes called the energy dimension. We include a brief argument in our setting for completeness.
@@
Let $\mu$ be a finite Borel measure on $M$ with correlation dimension $\dim_{C}(\mu)=s$ for some $s\in(0,n)$, and let $u\in(0,s)$. Then $I_{u}(\mu)<\infty$.
@@
$I_{u}(\mu)=\int_{0}^{\operatorname{diam}(M)}ur^{-u-1}F(r)\,dr+\mu(M)^{2}\,\operatorname{diam}(M)^{-u}.$
Fix $\varepsilon>0$ so that $u<s-\varepsilon$. Since $\dim_{C}(\mu)=s$, there exist constants $C>0$ and $r_{0}>0$ such that
$F(r)\leq Cr^{s-\varepsilon}\qquad(0<r\leq r_{0}).$
@@
$u\int_{0}^{r_{0}}r^{-u-1}F(r)\,dr\leq uC\int_{0}^{r_{0}}r^{s-\varepsilon-u-1}\,dr<\infty,$
since $s-\varepsilon-u>0$. On $r_{0},\operatorname{diam}(M)]$,
@@
Theorem 1.6 shows that for each fixed $u<s$ one can control the growth of the Kuznecov sum by $\lambda^{n-u}$ under the mild hypothesis $I_{u}(\mu)<\infty$. Our main theorem, Theorem 1.3, goes further by identifying the sharp exponent nsn-s and the exact leading constant in the asymptotic (1.7), under the additional assumption of the averaged $s$-density condition. As discussed in the introduction, the methods of [HR03] treat $u$ as a fixed parameter and do not provide the uniform control in $u$ near $s$ that is needed to pass to the endpoint $u=s$. This is why we turn to a more refined analysis based on the heat kernel and Tauberian theory.
5. Sharpness and examples
In this section, we discuss the sharpness of Theorem 1.3 and provide a few examples. We explain the necessity of the averaged $s$-density condition (1.5) for obtaining a genuine one-term asymptotic, and we show that the remainder $o(\lambda^{n-s})$ in (1.7) is essentially optimal at this level of generality. In particular, it cannot be improved uniformly to a power-saving error term, even over natural regular subclasses of measures. In the final part of this section we present an explicit self-similar fractal measure whose distance distribution $F(r)$ admits an averaged $s$-density constant, and hence satisfies (1.5).